One Leg Of A Right Triangle Is 21 Inches Longer Than The Smaller Leg, And The Hypotenuse Is 24 Inches Longer Than The Smaller Leg. Find The Lengths Of The Sides Of The Triangle.
Answer by mananth(16949)

One Leg Of A Right Triangle Is 21 Inches Longer Than The Smaller Leg, And The Hypotenuse Is 24 Inches Longer Than The Smaller Leg. Find The Lengths Of The Sides Of The Triangle.
...
smaller leg = x
longre leg = x+21
hypotenuse = x+24
...
Hyp^2 = leg1^2 + leg2^2
..
(x+24)^2=(x+21)^2+x^2
x^2+48x+576=x^2+42x+441+x^2
re write the equation
x^2+x^2-x^2-48x+42x-576+441=0
x^2-6x-135=0
x^2-15x+9x-135=0
x(x-15)+9(x-15)=0
(x-15)(x+9)=0
x= 15 ignore -9
...
The sides are 15, 36,39 inches
...
m.ananth@hotmail.ca